Two charges q1 = 4microC and q2 = -7 microC by 3.5m and 1.5m. Calculate the electric potential at point B having a distance of 2m.


1
Expert's answer
2022-02-13T15:22:35-0500

Assuming that the charges lie on the x-axis, and the distances given are from origin,


Potential at B due to q1


"V_1=k\frac{q}{r}"


"r_1= 3.5-2= 1.5m"

"V_1=\frac{9×10^9×4×10^{-6}}{1.5}=24000V"


"r_2= 2-1.5=0.5m"


"V_2 =\frac{9×10^9×-7×10^{-6}}{0.5}=-126000V"


Electric potential at B

"=-126000+24000=-102000V"